Weston Products manufactures an industrial cleaning compound that goes through three processing departments—Grinding, Mixing, and Cooking. All raw materials are introduced at the start of work in the Grinding Department. The Work in Process T-account for the Grinding Department for May is given below:

Work in Process—Grinding Department
Inventory, May 1 171,360 Completed and transferred
to the Mixing Department
?
Materials 405,020
Conversion 175,260
Inventory, May 31 ?

The May 1 work in process inventory consisted of 102,000 pounds with $116,280 in materials cost and $55,080 in conversion cost. The May 1 work in process inventory was 100% complete with respect to materials and 30% complete with respect to conversion. During May, 299,000 pounds were started into production. The May 31 inventory consisted of 130,000 pounds that were 100% complete with respect to materials and 60% complete with respect to conversion. The company uses the weighted-average method in its process costing system.

Required:

1. Compute the Grinding Department's equivalent units of production for materials and conversion in May.

2. Compute the Grinding Department's costs per equivalent unit for materials and conversion for May.

3. Compute the Grinding Department's cost of ending work in process inventory for materials, conversion, and in total for May.

4. Compute the Grinding Department's cost of units transferred out to the Mixing Department for materials, conversion, and in total for May.
Answer #1

Calculation of Physical Units

Beginning WIP + Started During the period = Completed & transferred + Ending WIP

Beginning WIP = 102,000

Started During the period = 299,000

Ending WIP = 130,000

Completed & Transferred =?

Substituting values to the equation

102,000 + 299,000 = Completed & Transferred + 130,000

401,000 = Completed & Transferred + 130,000

Completed & Transferred = 401,000 - 130,000

= 271,000

1)

Equivalent units for Materials :-

From Completed = 271,000

From Ending WIP(100%

Completed) (130,000*100%)=130,000

Total units = 401,000

  Equivalent units for Conversion Cost:-

From Completed = 271,000

From Ending WIP (60%

Completed) (130,000*60%) = 78,000

Total units = 349,000

Here Company uses Weighed average method so we not need to consider beginning WIP units and their percentage of completion. They are considered in FIFO method.

2)

Cost per Equivalent units for Materials :-

Cost from Beginning WIP = $116,280

Cost Added during the

Period = $405,020

Total Cost = $521,300

So equivalent units for material =

$521,300/401,000

= $1.3/Equivalent units

Cost per Equivalent units for Conversion Cost:-

Cost From Beginning WIP = $55,080

Cost Added During the

Period = $175,260

Total cost = $230,340

So equivalent units for conversion cost

= $230,340/349,000

= $.66/Equivalent units

As the company uses weighed average method we already included beginning units as fully in our calculation of equivalent units so we also have to include the cost of beginning WIP to calculate cost per equivalent units.

3)

  Cost of Ending WIP :-

To calculate ending WIP cost we have to just multiple cost per equivalent * ending WIP equivalent units

Materials = $1.3 * 130,000 = $169,000

Convention Cost = $.66*78,000 = $51,480

Total Ending WIP cost = $220,480

4)

Cost of units transferred out to Mixing Department :-

To calculate the cost of units transferred out we have to just multiple cost per equivalent * units completed & transferred out.

Material = $1.3*271,000 = $352,300

Conversion Cost = $.66*271,000 = $178,860

Total Cost transferred out = $531,160